Shift report 20011003 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 03 2001 - 08:00:10 EDT

  • Next message: Michael Murray: "Re: dndeta figures for impacients"

    ------------------------------------------------------------
    Date: Wed Oct  3 08:00:10 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: 
    Log book#9 pages 80 - 83
    
    Shift Summary:
    ------------------------------------------------------------
    1. Ramping attempts in the beginning of the shift, leading to 
       RHIC power supply front-end computer problem. A specialist had
       to be called for. After this a/the foil in the Tandem had to
       be changed. Therefore no beam by 06:00 when beam studies started.
       Also no beam by the end of shift.
    2. Two sets of pedestal and DVM runs.
    3. Since there was no beam the high background noted in evening
       shift could not be investigated further.
       TPM2 has got another set of 16 pads with wide pedestals. It is
       unknown to me why it occurs. It looks like a "hickup" somewhere in
       the read out of the pads. The 16 pads corresponds to one chip on
       a FEE card. (Some plots are put on the system status page, 
       http://pii3.brahms.bnl.gov/~operator/opdocs/SystemStatus.html)
       T1 has had some troublesome pads on row 8 for some time. The 
       pedestals for these pads have got worse, making it difficult to 
       subtract during data taking.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 03 2001 - 08:01:12 EDT